Obtain the two regression equation


X:10,12,14,16,18,20,22,24


Y:14,18,16,22,26,28,27,30 in mathematical foundations of computer science


1
Expert's answer
2022-01-11T19:18:58-0500

1) Regression equation ( Y on X):

Regression equation is written as:

"Y=a+bX"

Where;

"b=\frac{n\Sigma XY-\Sigma X \Sigma Y}{n\Sigma X^2-(\Sigma X)^2}"


"a=\frac{\Sigma Y-b\Sigma X}{n}"


Following table shows the calculation:




"b=\frac{(8\times3274)-(136\times181)}{(8\times2480)-(136)^2}=1.173"


"a=\frac{181-(1.173\times136)}{8}=2.690"


Then, regression equation is:


"Y=2.690+1.173X---------(1)"


2) Regression equation ( X on Y):

Regression equation is written as:

"X=a+bY"

Where;

"b=\frac{n\Sigma XY-\Sigma X \Sigma Y}{n\Sigma Y^2-(\Sigma Y)^2}"


"b=\frac{(8\times3274)-(136\times181)}{(8\times4349)-(181)^2}=0.776"


"a=\frac{\Sigma X-b\Sigma Y}{n}"


"a=\frac{136-(0.776\times181)}{8}=-0.556"


Then, regression equation is:


"X=-0.556+0.776Y---------(2)"